Factor completely: 3m^2 - 3n^2
Answer by ewatrrr(24785)

 
Hi,
Factor completely:
3m^2 - 3n^2 = 3(m^2-n^2) = 3(m-n)(m+n)
F First terms m^2
O Outside terms +mn
I Inside terms -mn
L Last terms -n^2